# IEEE-conference-template-062824.bib

### v1

@@ Abstract
Index Terms:
Object-oriented programming, Numerical algorithms, Parallel algorithms, Astronomy, Physics.
@@ I Introduction to CMB map-making
Cosmic inflation provides an elegant solution to a number of open problems in cosmology, and its several predictions have been tested as well. One distinct prediction of inflation is the production of stochastic gravitational waves (GW) that has not been observed yet. The B-mode polarization of cosmic microwave background (CMB) from recombination is predicted to carry the imprints of these stochastic GW from the early Universe [cmb1, cmb2, cmb3]. A number of space and ground-based experiments are therefore planned for the upcoming decade to probe the B-mode polarization of CMB [experiment1, experiment2, experiment3]. These experiments will employ thousands of detectors sampling the signal at very high frequencies, with data acquisition up to several hundred terabytes. Before proceeding with scientific analysis, these data must be reduced to sky maps in order to produce the intended scientific outcome of the experiment.
@@ II Data model and formalism
The matrix equation (3) is a single equation with two known quantities - $d$ and $P$ and two unknown quantities - $s$ and $n$. Our aim is to solve the matrix equation for the sky vector $s$. Assuming the noise vector to be Gaussian and stationary, we can write the linear solution for the sky vector $s$ as [ml1, ml2]:
@@
where $W$ is any positive-definite symmetric matrix compatible with $P$. For the noise covariance given by $N$, setting $W=N^{-1}$ in equation (4) will provide a minimum-variance estimate or the Generalized Least Square solution (GLS) [gls1] for the sky vector:
@@
Implementations of GLS written in the past have been instrumental in map-making for various ballon-based CMB experiments and space missions like WMAP [wmap] and Planck [planck]. Such implementations were successfully able to leverage the computational facilities to handle the data collected by the previous generation experiments. Though these implementations are still relevant, based on the projected data acquisition size, we estimate that they will require somewhere around ${\sim}200,000$ (for a typical space-based mission) to ${\sim}50,000,000$ (for a typical ground-based experiment) CPU hours to perform the map-making on data coming from future CMB experiments. If we include the map-making exercise that is performed during the development phase of these experiments, the computational cost for mission development will only increase. In light of future CMB experiments and the exponential growth in computational capabilities in the past decade, it is important that we revisit the map-making implementations and come up with the solution that can serve the computational demand of future experiments while taking full advantage of the new distributed and accelerated computing facilities. Our map-making framework BrahMap is one such initiative. With BrahMap we strive to provide a user-friendly map-making framework that can be used with a variety of computational facilities, including personal laptops, supercomputing clusters as well as multiple GPUs. In the rest of this paper, we will discuss the implementation details, features, performance scaling, and future scope of BrahMap.
@@ III-A Programming implementation
The programming implementation of BrahMap is based on COSMOMAP2111https://github.com/giuspugl/COSMOMAP2. BrahMap provides linear operator classes for the pointing matrix (and its transpose), block-diagonal preconditioner (see equation (7)), and several kinds of noise covariance matrix. We define these linear operators with an adopted version of LinearOperator class of SciPy [scipy] sparse linear algebra library (scipy.sparse.linalg [^2] https://docs.scipy.org/doc/scipy/reference/sparse.linalg.html). Linear operators defined with LinearOperator support mathematical operations such as addition, subtraction, multiplication, exponentiation, and inverse implicitly. In addition, they can also be used in the functions defined in scipy.sparse.linalg for matrix norms, solving linear problems, eigenvalue problems, singular value problems, and complete/incomplete LU factorization out-of-the-box. Most importantly, the LinearOperator class allows users to define their own linear operators independently that can integrate with the other operators seamlessly.
@@
    Python bindings for C++ routines: The Python interface of BrahMap and C++ computing backend are not naturally interoperable. One needs to bind the interface with the C++ function to ensure correct type conversion from one language to another. Among various options available for binding, we opted for pybind11 [pybind11]. pybind11 is a header-only library that exposes Python data types to C++ and vice versa. It offers native integration with NumPy [numpy] and C++ standard template library (STL) data structures. To prevent any situation of dangling pointers and memory leakage, we opted to allocate (and de-allocate) all the arrays in the Python side and pass them to C++ whenever they were required in computations.
@@ IV Future extensions to BrahMap
BrahMap is publicly available on Github [^7] https://github.com/anand-avinash/BrahMap with a detailed documentation [^8] https://anand-avinash.github.io/BrahMap. Using the features currently provided by BrahMap, it can be used to perform map-making on TODs for multiple detectors containing white noise. In the TODs coming from real experiment, the noise contribution is, however, complicated, and it also includes instrumental systematics. In order to perform map-making for such a dataset, it becomes important to introduce filtering templates and improved noise covariance models. Filtering template operators are usually sparse; they can be easily defined using LinearOperator class. Noise covariance, on the other hand, is dense. Given the enormous size of the noise covariance, it is impossible to load the entire dense matrix into memory. In such a case, we use some well-structured matrices, very close to the actual noise covariance, to obtain the maximum-likelihood maps using equation (4). A preferred choice for such structured matrix is the circulant matrix. Map-making with circulant noise covariance has been discussed elsewhere in the literature [mm1, mm2, mm3]. Since a circulant matrix can be diagonalized with the Fourier transform of its first row, we can compute matrix-vector multiplication with its inverse in just $O(n\log n)$ floating-point operations. Our work on circulant noise covariance operator (and its inverse) is in progress, and we plan to include its implementation in our next release.
@@ Acknowledgment
This work is supported by Italian Research Center on High Performance Computing, Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations). A.A. is also supported by the InDark and LiteBIRD INFN projects. We thank INFN LiteBIRD project for granting access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy) and the LEONARDO consortium.
[…]

### v2

# BrahMap: A scalable and modular map-making framework for the CMB experiments

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI: 10.1109/PDP66500.2025.00060
@@ I Introduction to CMB map-making
Cosmic inflation provides an elegant solution to a number of open problems in cosmology, and its several predictions have been tested as well. One distinct prediction of inflation is the production of stochastic gravitational waves (GW) that has not been observed yet. The B-mode polarization of cosmic microwave background (CMB) from recombination is predicted to carry the imprints of these stochastic GW from the early Universe [1, 2, 3]. A number of space and ground-based experiments are therefore planned for the upcoming decade to probe the B-mode polarization of CMB [4, 5, 6]. These experiments will employ thousands of detectors sampling the signal at very high frequencies, with data acquisition up to several hundred terabytes. Before proceeding with scientific analysis, these data must be reduced to sky maps in order to produce the intended scientific outcome of the experiment.
@@ II Data model and formalism
The matrix equation (3) is a single equation with two known quantities - $d$ and $P$ - and two unknown quantities - $s$ and $n$. Our aim is to solve the matrix equation for the sky vector $s$. Assuming the noise vector to be Gaussian and stationary, we can write the linear solution for the sky vector $s$ as [7, 8]:
@@
where $W$ is any positive-definite symmetric matrix compatible with $P$. For the noise covariance given by $N$, setting $W=N^{-1}$ in equation (4) will provide a minimum-variance estimate or the Generalized Least Square solution (GLS) [9] for the sky vector:
@@
Implementations of GLS written in the past have been instrumental in map-making for various ballon-based CMB experiments and space missions like WMAP [10] and Planck [11]. Such implementations were successfully able to leverage the computational facilities to handle the data collected by the previous generation experiments. Though these implementations are still relevant, based on the projected data acquisition size, we estimate that they will require somewhere around ${\sim}200,000$ (for a typical space-based mission) to ${\sim}50,000,000$ (for a typical ground-based experiment) CPU hours to perform the map-making on data coming from future CMB experiments. If we include the map-making exercise that is performed during the development phase of these experiments, the computational cost for mission development will only increase. In light of future CMB experiments and the exponential growth in computational capabilities in the past decade, it is important that we revisit the map-making implementations and come up with the solution that can serve the computational demand of future experiments while taking full advantage of the new distributed and accelerated computing facilities. Our map-making framework BrahMap is one such initiative. With BrahMap we strive to provide a user-friendly map-making framework that can be used with a variety of computational facilities, including personal laptops, supercomputing clusters as well as multiple GPUs. In the rest of this paper, we will discuss the implementation details, features, performance scaling, and future scope of BrahMap.
@@ III-A Programming implementation
The programming implementation of BrahMap is based on COSMOMAP2111https://github.com/giuspugl/COSMOMAP2. BrahMap provides linear operator classes for the pointing matrix (and its transpose), block-diagonal preconditioner (see equation (7)), and several kinds of noise covariance matrix. We define these linear operators with an adopted version of LinearOperator class of SciPy [12] sparse linear algebra library (scipy.sparse.linalg [^2] https://docs.scipy.org/doc/scipy/reference/sparse.linalg.html). Linear operators defined with LinearOperator support mathematical operations such as addition, subtraction, multiplication, exponentiation, and inverse implicitly. In addition, they can also be used in the functions defined in scipy.sparse.linalg for matrix norms, solving linear problems, eigenvalue problems, singular value problems, and complete/incomplete LU factorization out-of-the-box. Most importantly, the LinearOperator class allows users to define their own linear operators independently that can integrate with the other operators seamlessly.
@@
    Python bindings for C++ routines: The Python interface of BrahMap and C++ computing backend are not naturally interoperable. One needs to bind the interface with the C++ function to ensure correct type conversion from one language to another. Among various options available for binding, we opted for pybind11 [13]. pybind11 is a header-only library that exposes Python data types to C++ and vice versa. It offers native integration with NumPy [14] and C++ standard template library (STL) data structures. To prevent any situation of dangling pointers and memory leakage, we opted to allocate (and de-allocate) all the arrays in the Python side and pass them to C++ whenever they were required in computations.
@@ IV Future extensions to BrahMap
BrahMap is publicly available on Github [^7] https://github.com/anand-avinash/BrahMap with a detailed documentation [^8] https://anand-avinash.github.io/BrahMap. Using the features currently provided by BrahMap, it can be used to perform map-making on TODs for multiple detectors containing white noise. In the TODs coming from real experiment, the noise contribution is, however, complicated, and it also includes instrumental systematics. In order to perform map-making for such a dataset, it becomes important to introduce filtering templates and improved noise covariance models. Filtering template operators are usually sparse; they can be easily defined using LinearOperator class. Noise covariance, on the other hand, is dense. Given the enormous size of the noise covariance, it is impossible to load the entire dense matrix into memory. In such a case, we use some well-structured matrices, very close to the actual noise covariance, to obtain the maximum-likelihood maps using equation (4). A preferred choice for such structured matrix is the circulant matrix. Map-making with circulant noise covariance has been discussed elsewhere in the literature [15, 16, 17]. Since a circulant matrix can be diagonalized with the Fourier transform of its first row, we can compute matrix-vector multiplication with its inverse in just $O(n\log n)$ floating-point operations. Our work on circulant noise covariance operator (and its inverse) is in progress, and we plan to include its implementation in our next release.
@@ Acknowledgment
This work is supported by Italian Research Center on High Performance Computing, Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations). This work has also received funding by the European Union’s Horizon 2020 research and innovation program under grant agreement no. 101007633 CMB-Inflate. A.A. is also supported by the InDark and LiteBIRD INFN projects. We thank INFN LiteBIRD project for granting access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy) and the LEONARDO consortium.